The height of a triangle is 18.6 meters. Its area is 46.6 square meters. Find the length of the triangle's base

Answer by cleomenius(959)

Area = 1/2h*b
46.6 = 1/2 18.6* b
93.3 = 18.6*b
5 meters = base.
Cleomenius.